OpenStudy (anonymous):

A forest ranger spots a fire from a 28 foot tower. The angle of depression from the tower to the fire is 11 degrees. To the nearest foot how far is the fire from the base of the tower? Show steps

OpenStudy (anonymous):

You had it right the first time

OpenStudy (anonymous):

Just use tantheta

OpenStudy (anonymous):

using trigonometric ratios. tan 11=28/x

OpenStudy (anonymous):

I thought we were using sin ??

OpenStudy (anonymous):

therefore x = 28/tan 11

OpenStudy (anonymous):

tantheta = opp/adj, as your first picture depicted.

OpenStudy (anonymous):

I know what that means ^ But can you explain why x is adjacent or hypotenuse someone ?

OpenStudy (anonymous):